In a 2 digit no. , the tens digit is 3 times the units digit . When the no. is decreased by 54 the digits inter_changr . Find the no. ? ANS- Let no.s be x & y , 10x=y , y=2x Is it like this . THANKS
Answer by checkley71(8403)

30X+X-54=10X+3X OR 31X-13X=54 OR 18X=54 OR X=3
PROOF 30*3+3-54=10*3+3*3 OR 90+3-54=30+9 OR 93-54=39 OR 39=39